OpenStudy (anonymous):

7x+5y=-2 -6x+y=44 solve using substitution method

OpenStudy (watchmath):

Use the second equation and solve it for \(y\) to get \(y=44+6x\). Substitute this to the first equation to get \(7x+5(44+6x)=-2\) \(7x+30x+220=-2\) \(37x=-222\) \(x=-6\) It follows \(y=44+6(-6)=6\).

OpenStudy (anonymous):

would this be an ordered pair?

OpenStudy (watchmath):

There was a mistake \(y=44-36=8\) :) So the order pair is \((-6,8)\)